Fermi National Laboratory


2003 Accelerator Update Archive

December 29-January 5
- During this period of time Operations established six stores, which delivered luminosity for approximately 95 hours and 40 minutes.
- Booster accessed twice, first to repair a small LCW leak, and second to repair a RF cavity.
- Pbar accessed to repair a Septa.

December 22-December 29
- Tevatron vacuum leak in sector B1 requires cryo system warmup
- Antiproton Source Debuncher injection septum replaced due to vacuum leak
- Linac Klystron (KRF1) tube replaced several times
- Vacuum/cryo leak found on B14 spool piece

December 19-December 22
- During this period Operations established one store. That store plus another provided ten hours and twenty-one minutes of luminosity.
- A Linac quadrupole magnet ground faulted.
- On 12/19/03, the Fermi accelerators set a record for integrated luminosity of 2145 nb-1.
- Fermi set a record luminosity of 51.79E30
- A bad breaker at the Tevatron A0 Kicker building caused a six-house quench.
- A vacuum leak at Tevatron sector B14 or 15 required the cryo warm-up of B sector.
- The Tevatron will be down for about two weeks.

December 17-December 19
- Operations established two stores during this period that delivered 38 hours and 5 minutes of luminosity.
- Recycler shoot antiprotons to the Main Injector.

December 15-December 17
- Antiproton source broke stacking record of 236 mA with 245.017 mA
- Tevatron repairs continued

December 12-December 15
- Conning Towers check out completed
- Tevatron start up began with studies
- First store estimated at 48 hours away

December 10-December 12
- Conning Tower replaced
- Worked on loss problems in Booster and MiniBooNE

December 8-December 10
- Studies and maintenance work due to Tevatron conning tower replacement

December 5-December 8
- There were no stores established during this period. The one store that already existed at the beginning of this period provided 2 hours and 36 minutes of luminosity.
- Switchyard receives permission to receive beam.
- The Tevatron suffers a sixteen-house quench, which sets an unfortunate record.
- Cryogenic problems shuts down the Tevatron for a week.

December 3-December 5
- Operations established two store during this period, which offered 16 hours and twenty-four minutes of luminosity.
- Recycler Pbar transfers suffer due to a RF mismatch
- RF cavity in Pbar causes problems

December 1-December 3
- The integrated luminosity for last week was 3.3 pb-1
- TeV experts work on quench-at-abort problems
- Booster achieves it first 5E12 pulse of beam

November 26-December 1
- Meson feeder 30 was back online Wednesday morning
- Main Injector suffers serious LCW leaks early Thanksgiving day
- MiniBooNE has Horn problem and then airflow problems on Saturday
- Linac has debuncher problems
- During this time period Operations established four stores. With this and an existing store, the accelerators delivered approximately 101 hours and 29 minutes of luminosity.

November 24-November 26
- Meson feeder 31 faults again
- A1 and C4 wet engines require rebuilds
- TeV establishes first official store

November 21-November 24
- Operators forced to switch Preacc operations to I- Source due to a rain leak
- Operations and TeV experts establish a study-store with a 0.9E31 luminosity
- Meson feeder 30 faulted and caused CHL power glitch

November 19-November 21
- All machines conduct final studies
- The Antiproton Source successfully shot antiprotons to the Main Injector
- Collider physics will hopefully resume this weekend

November 17-November 19
- Problem with the Tevatron C2 dump delays studies six hours
- Tevatron beam accelerated to flattop
- The H- Source suffers from rainfall

November 14-November 17
- Linac Klystron RF station 4 failed and needed replacing
- Beam established to Booster, Main Injector, and the Antiproton Source
- CDF powers their solenoid
- MiniBooNE begins receiving beam
- The Antiproton Source began stacking
- Tevatron circulating beam established
- Recycler circulating beam established

November 12-November 14
- ComEd and Fermi high voltage static line problem cause two site wide power outages.
- Shutdown ends.

November 10-November 12
- Cable pull between TeV sectors A1 and A2 finished
- Future Plans
- The Fermi Accelerator Complex

November 7-November 10
- Operations received permission to run beam in the Main Injector
- Feeder 45 trip caused problems for the Tevatron, Linac, Booster, and CHL
- Experts accelerated MI beam to 150 GeV
- Future Plans

November 5-November 7
- The Tevatron ramped to 1005 GeV
- A power glitch trips off CHL's nitrogen plant
- Tevatron Status
- A Basic Refrigeration System

November 3-November 5
- Tevatron cold by Tuesday morning
- Operations given permission to power the TeV
- Accelerator Status

October 31-November 3
- Booster suffers LCW leak
- Beam established to MiniBooNE
- Feeder 31 trips off parts of Meson
- The Meson area

October 29-October 31
- Booster leaks resolved and Booster circulates beam
- MiniBooNE power supplies tested
- Tevatron Status
- Recycler Status

October 27-October 29
- Booster Activities

October 24-October 27
- Electricians continue their TeV service building maintenance
- Operators successfully interlock and highpot TeV sector F, MI, and Booster
- Operations run devices in Booster and Main Injector
- Main Injector quadrupole found with internal ground fault
- Antiproton Source ground fault discovered and repaired

October 22-October 24
- Tevatron bake outs finished
- Linac switched to the H- Source
- Tevatron Status
- Japanese Pots
- Recycler Status

October 20-October 22
- Linac technicians replace two RF power amplifiers
- Booster RF stations tested
- The Cockcroft-Walton Accelerator

October 17-October 20
- Linac and Booster suffers from a few turn on problems
- Main Injector has planned power outage

October 15-October 17
- Linac established beam to its dump on Thursday
- Booster began running all its RF stations 60%
- Recycler Status
- Tevatron Status

October 13-October 15
- Linac startup has begun
- Recycler: Electron Cooling

October 10-October 13
- ComEd's planned power outages
- The MiniBooNE experiment

October 8-October 10
- Small leaks develop as LCW systems resume operation
- Antiproton Source motorized quadrupoles
- Tevatron work status

October 6-October 8
- The Main Injector water system turned on
- Pictures of the Antiproton Source motorized quadrupole magnets
- Installation of the Tevatron HLS will finish this week

October 3-October 6
- CUB turned on the 55-LCW system
- Shellfish mussel in on the Tevatron

October 1-October 3
- Tevatron Status
- Switchyard Status
- Linac Status
- Recycler Status

September 29-October 1
-High voltage maintenance work put CDF in the dark
-Pond pump line develops leak at F2
-Linac and Booster pictures: Linac water system, Booster collimator, and Linac Lambertson

September 26-September 29
- A weekend full of power outages
- Tevatron and Main Injector heat exchanger cleaning

September 24-September 26
- A Recycler problem reported on Monday was not due to the Main Injector, but due to vacuum readback problems.
- Central Helium Liquefier building and all the Tevatron service buildings suffered from a power outage at 4 AM on Friday morning.
- Tevatron Status
- Booster Status

September 22-September 24
- Main Injector vacuum causes problems for the Recycler
- TeV Hydrostatic Level System (HLS) installation

September 19-September 22
- On Saturday technicians powered down the Education Center, Linac, and MiniBooNE for maintenance.
- Cryogenic dipole problem discovered in the Tevatron.

September 17-September 19
- F0 Lambertson gets new liner
- List of Antiproton Source shutdown work

September 15-September 17
On Monday, the Beams Division Head revoked permission to operate all of the accelerators except for the Preaccelerator. Before an accelerator can run again each accelerator will have to perform the required safety checks and complete a restart form, which the Division Head must approve. Around 9 AM the power for the MI-8 transfer line from Booster to the Main Injector (MI) was switched to a generator.

On Tuesday morning, the MI LCW system was shut down to begin cleaning the heat exchangers. Around 6 PM a Collider Detector Facility experimenter reported a Kautzky valve that was leaking nitrogen. Cryo technicians came in and repaired the valve.

September 12-September 15
- The bake-out of the first Recycler vacuum sections is in progress.
- The location of the coils in 20 Tevatron magnets have been corrected with shims.
- All core drilling for the monuments of the new TeV Net, a new surveying network, is finished.

September 10-September 12
- The first week of the shutdown is over. Among other things, work was done on Lambertson liners, Collimator and Conning tower.

September 8-September 10
- A power outage at the Main Injector 20 area caused problems and delayed some shutdown work.

September 5-September 8
- During this seventy-two hour period, Operations established one store. That store, plus an already existing one provided twenty-two hours of luminosity to the experiments.
- The Tevatron suffered a quench while conducting a dry squeeze.
- A ComEd power glitch hits the Kautz Road substation.
- The seven to ten week shutdown began at 8 AM today, September 8, 2003

September 3-September 5
- During this forty-eight hour period, Operations established two stores. They delivered approximately twenty-eight hours and fourteen minutes of luminosity to the experiments.
- The TeV suffered a quench due to a problem with dampers.
- A Pbar Debuncher device (D:ISEPV) temporarily halts stacking.

September 1-September 3
- Operations established one store during this forty-eight period. This store, along with one other provided approximately twenty-one hours and thirty-one minutes of luminosity to the experiments.
- A TeV store aborted due to a blown high voltage fuse at the F0 service building.
- Experts successfully completed seven 2.5MHz transfers of antiprotons from the Main Injector to the Tevatron.

August 29-September 1
- During this seventy-two hour period, Operations established two stores that delivered eighteen hours and fifty-one minutes of luminosity to the experiments.
- Two potential TeV stores lost due to quenches. Experts discover cogging problem.

August 27-August 29
- Operations established one store during this forty-eight hour period. Including an already established store, these two stores delivered approximately six hours and fifty-five minutes of luminosity to the experiments.
- Store 2956 quenched due to a bad VFC card.
- Thursday, August 28, was a maintenance day.

August 25-August 27
- Operations established two stores during this forty-eight hour period and delivered thirty-two hours and twenty minutes of luminosity to the experiments.
- A lightening strike caused the Tevatron to quench.
- Fermilab domestic water was off last night due to a leak. C0 and D0 domestic water still off.

August 22-August 25
- During this past seventy-two hour period, Operations established three stores. These stores delivered approximately sixty-one hours and twenty-eight minutes of luminosity to the experiments.
- A feeder 47 power glitch caused the Tevatron to quench.

August 20-August 22
- Operation established two stores during this period. Those stores delivered approximately thirty-hours and nine minutes of luminosity to the experiments.
- Fermilab reached its goal for this fiscal year of delivering a total integrated luminosity of 225 pb-1. at 6:53 AM on Friday, August 22.
- A ComEd power glitch caused the loss of the Pbar stack.

August 18-August 20
- Quenches trouble the Tevatron
- A TeV RF station's new power amplifier tube fails and is replaced
- Operations established three stores during this period. Those stores provided six hours and forty-nine minutes of luminosity to the experiments.

August 15-August 18
- Booster suffers vacuum leak that required access and repair. Many other areas took advantage of this downtime and performed maintenance.
- A Linac RF station, LRF2, was off for about three hours due to a bad relay.
- The Meson LCW system suffered a leak.
- Operations established three stores during this period. Including an already established store, the accelerators provided approximately forty-nine hours and thirty minutes of luminosity to the experiments.
- Fermilab has accumulated approximately 221.9 inverse pico barns. Our goal for this fiscal year was 225 inverse pico barns.

August 13-August 15
- Tevatron suffered compressor, RF and VFC problems.
- Operations established three stores during the last forty-eight hours.
- They delivered approximately nineteen hours and forty five minutes of luminosity to the experiments.

August 11-August 13
-Tevatron suffers multiple quenches.
- Operations established one stores during the last forty-eight hours.
- They delivered approximately ten hours and fifteen minutes of luminosity to the experiments.

August 8-August 11
- Tevatron store #2887 sets new record for initial luminosity.
-Tevatron also suffered two quenches, high losses, and RF problems.
- Operations established three stores during the last seventy-two hours. They delivered approximately twenty-eight hours and fifteen minutes of luminosity to theexperiments.

August 6-August 8
- Linac RF station 2 required a lot of expert attention
- An Antiproton Source pretarget beamstop required maintenance
- Operations established one store and monitored another during this forty-eight hour period. The stores delivered approximately eighteen hours and twenty-seven minutes of luminosity.

August 4-August 6
- The Tevatron suffers quenches, closure, orbit, and quadrupole troubles
- A Feynman Computer Center power outage caused network problems
- Operations established two stores during the last forty-eight hours. They delivered approximately twenty-three hours and twenty-one minutes of luminosity to the experiments.

August 1-August 4
- RF problems troubled nearly every machine
- Experts trace Linac instability to a faulty multiwire
- The Tevatron Electron Lens suffers an anode short
- Operations established one store during this period. That store plus one already started on Friday delivered approximately twenty-six hours and nine minutes of luminosity.

July 30-August 1
- ComEd admits causing two power glitches, which tripped off Booster power supplies for the previous morning
- Tevatron quenches take toll on HEP
- Operations monitored a store from Wednesday and established one store during this period. These two stores provided eleven hours and twenty-two minutes of luminosity.

July 28-July 30
- Cryo and vacuum repairs made late Monday evening and Tuesday owl shifts.
- Snake slows operator response to reset tripped TeV RF station.
- The integrated luminosity for last week was 9.7 pb-1.
- There were two stores established during this period. The Tevatron delivered luminosity for approximately thirty-two hours and fifty-one minutes.

July 25-July 28
- New initial luminosity record of 45.34e30!

July 23-July 25
- A bad temperature sensor may be the cause of the many Booster power supply (VCB1) trips. This will require an access to repair.
- The Tevatron lost two stores due to a quench and a vacuum burst.
- TeV store 2817 began with an initial luminosity of 42.64E30. The highest luminosity for this run was established on 5/17/03 with 44.89E30.
- Operations established two stores during this time period. There was a total of 34 hours and 18 minutes of luminosity delivered.

July 21-July 23
- The integrated luminosity for last week was 5106.9 nb-1
- Cryo experts installed new software that should halt the valve excursion problems
- There were two stores established during this time period. They delivered luminosity to the experiments for thirty-five hours and one minute.

July 18-July 21
- Pbar has a kicker power amplifier tube that is failing. The experts think it will last until the shutdown.
- Switchyard experts establish beam to the Meson Target Train.
- A lightening strike early Monday morning drastically reduces store 2805's luminosity and causes many devices to trip off.
- There were four stores established during this time period for a total duration of forty-eight hours and twenty-seven minutes.

July 16-July 18
- Late Thursday afternoon a lightning induced power glitch caused a store to abort and many devices to trip off. However, the Recycler maintained its store of protons.
- Operations established two stores during this period. The total duration for both stores was nine hours and thirteen minutes

July 14-July 16
- A TeV quadrupole power supply failed late Monday night. EE support people worked throughout the night until 6 AM Tuesday morning to get the supply back into operation.
- A lightning strike at 4:35 AM on Tuesday morning played havoc with the TeV cryo system at C4
- Pbar suffered from emittance and LCW problems
- There was only one store during this troublesome period. It lasted for eighteen hours and 25 minutes.
- The integrated luminosity for last week was 8467.4 nb-1.

July 11-July 14
- Booster kicker problems require magnet change
- Tevatron RF problems cause drop in luminosity and loss of stores

July 9-July 11
- Preaccelerator experts switched from the H- Source to the I- Source
- The Tevatron suffered from various LCW problems
- The Tevatron cryo system also needed work

July 7-July 9
- Lightning induce power glitch causes store 2768 to abort.
- Cryo technicians repair wet engine at TeV sector C3.
- During this time period the complex had two good stores and started a third.

July 2-July 7
- As the Operations Department head stated, "This was the worst operational weekend I've seen in thirty years." What more needs to be said?

June 30-July 2
- The second highest weekly integrated luminosity was achieved last week with 8629.90 nb-1
- Store 2736 lost due to cryo valve controls problem
- H- Source has gas valve failure. I- Source started up

June 27-June 30
- Linac suffering quadrupole problems that disturbed the beam for Booster and MiniBooNE
- CDF and DZero accessed their collision halls

June 25-June 27
- Technicians replace the faulty MiniBooNE quadrupole power supply breaker

June 23-June 25
- The Collider Detector Facility requires access to fix power supply
- Tevatron experts investigating quenches at sector BA

June 18-June 20
- Switchyard proves that it can take beam simultaneously with the Tevatron operating at 980 GeV.
- The NuMI project has a power outage.
- The CDF Roman Pots suffer a control problem.

June 16-June 18
- Switchyard experts preparing to send beam to Meson

June 13-June 16
- The power glitch caused by a fault on Feeder 44 occurred Friday afternoon and caused some systems to trip. The power was rerouted and this cable will be repaired later.
- Over the weekend we did mostly tuning of the machines and put stores into the Tevatron.

June 11-June 13
- Booster and Linac problems halts beam delivery.
- An Accumulator device fails and causes the loss of the stack.

June 9-June 11
- A Switchyard expert sent beam to TeV sector F49 on the first pulse.

June 6-June 9
- On Friday an Antiproton Source Kicker problem halted stacking, and on Saturday a Linac microprocessor inhibit did the same. Both problems were fixed by the mere presence of an expert.
- A wet engine failed early Sunday morning and held off beam until noon.

June 4-June 6
- A TeV quench caused the loss of store 2644 on Wednesday.
- A Debuncher tunnel vacuum problem required a Thursday evening access to repair.

June 2-June 4
- A Linac RF station's power amplifier is failing (LRF1).
- The seven-week summer shutdown will begin on August 25th.

May 30-June 2
- High Voltage power pole replaced by ComEd.
- The Fermilab accelerator complex recovers from both planned and unplanned power outages.
- High Energy Physics (HEP) finally resumes early on Monday morning.

May 28-May 30
- Accelerator complex in shutdown
- Scheduled site wide power outage to switch power from one substation to another
- Unscheduled site wide power outage followed

May 23-May 28
- Quadrupole shunt SCR discovered off in AP3 line
- Booster expert discovered RF phase problems
- Planned power outage to replace a bad high voltage powerpole
- Accelerator complex will be in shutdown for three days

May 21-May 23
- Quench trips caused by a barrel spice failure
- Pbar experts successfully install and test new lithium lens

May 19-May 21
- New Main Injector record - high intensity of protons, 5.45E12, sent to the Antiproton Source (Pbar) target for a single pulse

May 16-May 19
- New integrated luminosity record with a level of 9159 nb-1
- New initial luminosity record of 44.895E30
- New antiproton stacking-rate record of 13.51mA in one hour

[an error occurred while processing this directive]

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

Return to Current Accelerator Update



last modified 1/5/2004   email Fermilab